OpenStudy (anonymous):

i know it is basic but it has been awhile 3(x-5) over 2 = 2(x+5) over 3

OpenStudy (amistre64):

clear the fractions by "cross multiplying"

OpenStudy (amistre64):

then go in for the win

OpenStudy (anonymous):

if you'd like to share your answer we'll tell you if your right

OpenStudy (anonymous):

do take the 2 and mulitiply the the other side ?

OpenStudy (amistre64):

yes, and the same with the 3

OpenStudy (anonymous):

idea is \[\frac{a}{b}=\frac{c}{d}\iff ad=bc\]

OpenStudy (amistre64):

then expand it by multiplying thru the (...)

OpenStudy (anonymous):

I have already took out the ()

OpenStudy (amistre64):

thats fine too :)

OpenStudy (amistre64):

once youve cleared the fractions; gather like terms

OpenStudy (anonymous):

do i multiply the 3x and2x?

OpenStudy (amistre64):

\[\frac{3(x-5)}{2} = \frac{2(x+5)}{3}\] \[\frac{3(x-5)*(3)(2)}{2} = \frac{2(x+5)*(3)(2)}{3}\] \[3(x-5)*(3) = 2(x+5)(2)\] \[9(x-5) = 4(x+5)\] \[9x-45 = 4x+20\] this is what I get to

OpenStudy (amistre64):

9x - 45 = 4x + 20 ; +45 +45 +45 ---------------- 9x = 4x + 65 ; -4x -4x -4x ---------------- 5x = 65 5x = 65

OpenStudy (amistre64):

any of that ring a bell :)

OpenStudy (anonymous):

yes thanks

OpenStudy (amistre64):

youre welcome :)

OpenStudy (anonymous):

\[3/4x\]

OpenStudy (anonymous):

3/4x - 1/2=1 how would i got about this one

OpenStudy (anonymous):

x=13 remove denominators by multiplying to the top of each line cancel them out and the rest is easy

OpenStudy (anonymous):

9x add 4x = 20 + 45

OpenStudy (anonymous):

i am on a different equation now